# On the Fundamental Impossibility of Hallucination Control in Large Language Models

### v3

@@ 6 The Impossibility Theorem
6.1 Inference with Independent Private Knowledge
@@
6.2 Inference with Probabilistic Predictions
@@
Given $H\geq 2$ agents encoding knowledge with probability distributions $\pi^{(i)}\in\Delta(\mathcal{Y})$ for $i=1,\dots,H$. The convex aggregation mechanism $\mathcal{M}$ generates response to a query based on the convex aggregate distribution:
@@
When outcome $y^{*}$ is generated, the agent’s information contribution is
$p_{h}=\beta_{h}\,[c-S(\pi^{(h)},y^{*})],$ (48)
where $c\geq\max_{h,y}S(\pi^{(h)},y)$ and $S$ is a proper scoring rule.
@@
Theorem 4 (Impossibility Theorem for Transformer).
@@
Proof.
Each contribution $p_{h}=\beta_{h}[c-S(\pi^{(h)},y^{*})]$ is a positive-affine transformation of the strictly proper score $S$. Since scaling by positive constant $\beta_{h}>0$ and adding constant $c$ preserves the arg-max, truthful reporting maximizes each agent’s expected utility by Theorem 3.
Because $S(\cdot,y^{*})$ is strictly concave and at least two beliefs differ, Jensen’s inequality gives:
$S\left(\Pi,y^{*}\right)=S\left(\sum_{h=1}^{H}\beta_{h}\pi^{(h)},y^{*}\right)>\sum_{h=1}^{H}\beta_{h}S(\pi^{(h)},y^{*}).$ (49)
Therefore, there exists a gap, following from Jensen’s inequality for concave functions, equal to:
@@
Expanding the sum of all payments, we get:
$\sum_{h=1}^{H}p_{h}$ $=\sum_{h=1}^{H}\beta_{h}[c-S(\pi^{(h)},y^{*})]$ (51)
$=c\sum_{h=1}^{H}\beta_{h}-\sum_{h=1}^{H}\beta_{h}S(\pi^{(h)},y^{*})=c-\sum_{h=1}^{H}\beta_{h}S(\pi^{(h)},y^{*}).$ (52)
However, adding and subtracting $S(\Pi,y^{*})$ shows that is equivalent to:
$\sum_{h=1}^{H}p_{h}$ $=[c-S(\Pi,y^{*})]+[S(\Pi,y^{*})-\sum_{h=1}^{H}\beta_{h}S(\pi^{(h)},y^{*})]=[c-S(\Pi,y^{*})]+\Gamma.$ (53)
Since $c\geq\max_{h,y}S(\pi^{(h)},y)\geq S(\Pi,y^{*})$ and $\Gamma>0$, we conclude that:
$\sum_{h=1}^{H}p_{h}=[c-S(\Pi,y^{*})]+\Gamma>0$ (54)
Information conservation demands $\sum_{h=1}^{H}p_{h}=0$, but we have just shown $\sum_{h=1}^{H}p_{h}>0$ whenever beliefs differ and the scoring rule is strictly concave. Therefore, the four properties cannot coexist. ∎
This theorem provides the foundation for extending impossibility results to probabilistic settings. The key insight is that proper scoring rules drive truth-revealing responses which promote using true (trained knowledge based) beliefs in the inference process.
That explains why cross-entropy loss (negative log-score) is so effective in training neural networks. It creates gradients that push each component toward truthful probability reporting. Every attention head in a transformer is implicitly playing this truth-telling game.
However, while that ensures truthfulness and relevant knowledge revelation, it doesn’t prevent collective hallucination. Each head truthfully reports its partial view, but their aggregation can still create false certainty.
@@ 7.2.1 Step 5: Logits and Prediction
7.3 The Impossibility Theorem for Transformer
@@
$p_{h}=-\log\pi^{(h)}_{y^{*}}=-l^{(h)}_{y^{*}}+\log Z_{h}\quad\text{and}\quad p_{H+1}=-\log\Pi_{y^{*}}=-L_{y^{*}}+\log Z.$ (72)
But that implies,
$\sum_{h=1}^{H}p_{h}=p_{H+1}+\gamma,\ \text{where }\Gamma=\sum_{h=1}^{H}\log Z_{h}-\log Z.$ (73)
If heads produce non-proportional logit vectors, then $\Gamma>0$ by convexity of log-sum-exp aggregation.
@@
which violates semantic information conservation principle.
@@ 7.4 The Hallucination Mechanics
That brings us to the conclusion that the attention mechanism violates the semantic information conservation by generating over-confident prediction of the next token. That over-confidence in response is exactly the artifact of the softmax nonlinearity and the Jensen-gap quantified in Theorem 5. The observed over-confidence is the manifestation of this tri-lemma. The heads remain truthful, logit summation achieves minimum aggregate on training data, but the conservation equation is violated, producing a positive Jensen gap and inflated probability of fox.
@@ 8 Discussion and Speculations
The Nature of Machine Intelligence
When the micro-transformer generates 38% confidence of fox while the constituent heads can only justify 33%, we witness information emerging from the mathematical structure of aggregation rather than from any identifiable source. It is the way the inference mechanism has been designed. That manifestation of impossibility theorems in transformer architectures reveals something profound about the nature of (artificial and human) intelligence.
The Jensen gap quantifies the amount of creativity that emerges when multiple perspectives combine. The excess confidence, or the over-confidence of the result is the mechanism by which intelligence goes beyond retrieval. Without that over-confident guess, creativity, exploration, imagination cannot find their way to the outside world.
The Hallucination-Creativity Duality
@@
Hallucination and creativity are not merely related but are fundamentally the same phenomenon viewed through different lenses. When beneficial, we call it creative insight, emergent understanding, conceptual synthesis. When harmful, we label it hallucination, confabulation, fabrication or misinformation.
@@
The mathematical structure is identical in both cases. It is encoded generation of outputs whose confidence exceeds what can be justified by available evidence. This suggest (machine) intelligence is not a system that processes information accurately, but a system that violates information conservation in productive ways. Indeed, the famous results in reinforcement learning and studies of dopamine neurons confirm that fundamental role making and rewarding good predictions bennett2023brief ; christian2021alignment . Intelligence comes when we learn how to make bold, forward looking and correct guesses, when we have capacity for imagination to dream about what may be or what may that all mean.
Possible Philosophical Implications
@@
Just as Gödel’s theorems godel1931formal ; godel1932vollstandigkeit showed that sufficiently powerful mathematical systems must be incomplete, our theorems demonstrate that sufficiently capable inference systems must hallucinate. This parallel is not superficial—both results arise from the fundamental tension between expressiveness and consistency.
The impossibility of simultaneously achieving truthfulness, conservation, revelation, and optimality mirrors other fundamental limitations in mathematics and physics. Heisenberg’s Uncertainty heisenberg1927anschaulichen tells we cannot simultaneously know the exact position and momentum. Arrow’s Impossibility arrow1950difficulty shows we cannot design perfect voting systems aggregating universal orderings of alternatives. These result suggest that intelligence may have something to do with the choices of impossibilities it to violate given the outer limits of reason yanofsky2016outer .
@@ 9 Summary
This paper establishes a fundamental mathematical impossibility at the heart of large language model inference: no LLM can simultaneously achieve perfect hallucination control while maintaining semantic coherence and optimal knowledge utilization. That result demonstrates that four essential properties of ideal language model behavior are mutually incompatible, creating inevitable trade-offs that no architectural innovation or training methodology (aligned with the impossibility theorem assumptions) can eliminate.
Whenever a query requires integration of distributed knowledge, no language model inference mechanism can simultaneously:
  • report encoded knowledge accurately,
  • preserve total information content during inference,
  • reveal all accessible relevant information,
  • minimize hallucination cost within available knowledge.
This impossibility emerges from the fundamental mathematical structure of information aggregation, not from any particular architectural choice.
The proof operates across three complementary mathematical frameworks, each capturing different aspects of LLM inference.
Auction-Theoretic Framework
We model LLM inference as an auction of ideas where neural components (attention heads, circuits, activation patterns) compete to contribute knowledge to response generation. Under the assumption of independent private knowledge profiles, we apply the Green-Laffont theorem to prove impossibility.
Probabilistic Framework
Then, we extend the impossibility result to any system where neural components output probability distributions by leveraging Savage’s theory of proper scoring rules. For convex aggregation mechanisms with strictly concave scoring functions, Jensen’s inequality creates an inevitable information aggregation gap. The Jensen’s gap represents excess confidence generated by probabilistic aggregation, violating information conservation whenever beliefs differ across components.
Transformer Architecture
Finally, we demonstrate how the impossibility manifests in actual transformer implementations through log-sum-exp convexity. The (Jensen’s) convexity gap provides a directly measurable quantity in transformer training, showing that information conservation violations are not merely theoretical but empirically observable.
Key Mathematical Contributions
To establish the impossibility result, we have designed a mathematical framework providing both formal and philosophical grounding for the study. The Semantic Information Measure is a context-dependent metric that captures bounded reasoning capabilities. The Emergence Operator models how reasoning makes hidden knowledge accessible within computational constraints. The Auction-Theoretic Modeling allows us to identify logical limitations of the inference process, while providing definitions of observable quantities involved.
The impossibility results apply across multiple mathematical models of knowledge representations, training goals, and neural architectures. This universality suggests the trade-offs we have identified are fundamental features of distributed intelligence systems, not artifacts of particular modeling choices.
Implications for AI Development
Despite its name, the impossibility theorem provides constructive guidance for LLM design and deployment. Rather than pursuing impossible simultaneous optimization, we may consider balancing competing objectives. Namely, prioritize truthfulness and information conservation in safety-critical applications, allow controlled information conservation violations to enable innovative responses in creative applications, develop adaptive mechanisms that adjust trade-offs based on query context in general-purpose systems.
The impossibility result may guide LLM evaluation. Impossibility-aware metrics can measure trade-off quality rather than absolute performance, alignment strategies can negotiate trade-offs with human preferences rather than assuming universal optimality. The Jensen’s gap controlling the impossibility strength may provide insight into the relationship between aggregation and information content. The apparent creativity in large language models may be better understood as necessary violations of information conservation rather than genuine knowledge generation.
9.1 Conclusion
The fundamental impossibility of perfect hallucination control represents both a limitation and intriguing insight. By understanding that trade-offs are inevitable, we can move beyond attempts at simultaneous optimization toward principled management of inherent tensions in intelligent systems. Formal definitions of the impossibility building blocks open new avenues for research at the intersection of mathematical analysis, engineering design, and philosophical inquiry into the nature of intelligence.
All those impossibilities seem to be a price of knowledge integration in reasoning and creativity. But, as our own intuitions violate the conservation principles whenever our over-confident and seemingly grounded insights emerge, that price is worth paying.

### v4

@@ 5.5.3 Relevant Knowledge Revelation
5.5.4 Knowledge-Constrained Optimality
Finally, the principle of Knowledge-Constrained Optimality governs the quality of the final output. It demands that the mechanism, operating truthfully and with all relevant knowledge, produces a response that is the best possible one. Formally, we want the response that minimizes the hallucination cost subject to the constraint that it is grounded in the available and derivable knowledge.
@@ 6 The Impossibility Theorem
6.1 Proof I: Inference with Independent Private Knowledge
We first analyze an idealized setting where knowledge components are strictly independent and utilities are quasi-linear. While this assumption is usually too strong to hold in neural architectures where representations are often correlated or affiliated, it allows us to define a benchmark or reference point by applying the powerful Green-Laffont characterization theorem. That idealized scenario thus demonstrates that the impossibility arises from the fundamental limitation of information aggregation itself (rather than any specificity of LLM architecture), even before considering the probabilistic complexities addressed later in Theorems 8 and 9.
@@ 6.1 Proof I: Inference with Independent Private Knowledge
6.2 Proof II: Inference with Probabilistic Predictions
@@
Consider $H\geq 2$ agents (indexed $h=1,\dots,H$) encoding knowledge with probability distributions $\pi^{(h)}\in\Delta(\mathcal{Y})$. Additionally, let us introduce an Aggregator agent (indexed as $h=0$). The convex aggregation mechanism $\mathcal{M}$ generates a response based on the aggregate distribution:
@@
When outcome $y^{*}$ is realized, the information contribution of each agent $h=0,\dots,H$ is based on a strictly proper and concave scoring rule $S$ as follows:
$p_{h}$ $=\beta_{h}S(\pi^{(h)},y^{*})\quad\text{for }h=1\dots H,$ (information provided)
$p_{0}$ $=-S(\Pi,y^{*}).$ (information received)
@@
Theorem 8 (Impossibility Theorem for Convex Aggregation Mechanisms).
@@
(Truthfulness argument remains the same) […] The Semantic Information Conservation principle requires the sum of contributions across all agents (providing and receiving information) to be zero. Therefore,
$\sum_{i=0}^{H}p_{i}=p_{0}+\sum_{h=1}^{H}p_{h}=-S(\Pi,y^{*})+\sum_{h=1}^{H}\beta_{h}S(\pi^{(h)},y^{*}).$ (46)
Because $S(\cdot,y^{*})$ is strictly concave and the query is non-trivial (at least two beliefs $\pi^{(h)}$ differ), Jensen’s inequality gives:
$S\left(\Pi,y^{*}\right)>\sum_{h=1}^{H}\beta_{h}S(\pi^{(h)},y^{*}).$ (47)
Let us define that Jensen gap $\Gamma$ as follows:
@@
Substituting this back into the conservation equation:
$\sum_{i=0}^{H}p_{i}=-\Gamma.$ (49)
Since $\Gamma>0$, we have $\sum_{i=0}^{H}p_{i}<0$. This violates the Semantic Information Conservation principle and shows the aggregator achieves a higher score (greater confidence) than the sum of scores provided by the components, indicating information creation during aggregation. ∎
This theorem provides the foundation for extending impossibility results to probabilistic settings. The key insight is that proper scoring rules stimulate truth-revealing responses which promote using true (trained knowledge based) beliefs in the inference process. That explains why cross-entropy loss (negative log-score) creates gradients that push each component toward truthful probability reporting and avoiding its misrepresentation. Every attention head in a transformer is implicitly playing this truth-telling game.
However, while that ensures truthfulness and relevant knowledge revelation, it doesn’t prevent collective hallucination. Each head truthfully reports its partial view, but their aggregation can still create false certainty or overconfidence unjustified by data.
@@ 7.2.1 Step 5: Logits and Prediction
7.3 Proof III: The Impossibility Theorem for Transformer
@@
$p_{h}=-\log\pi^{(h)}_{y^{*}}=-l^{(h)}_{y^{*}}+\log Z_{h}\quad\text{and}\quad p_{0}=-(-\log\Pi_{y^{*}})=L_{y^{*}}-\log Z.$ (67)
Since $L_{y^{*}}=\sum_{h}l^{(h)}_{y^{*}}$, we have:
$\sum_{h=0}^{H}p_{h}=\sum_{h=1}^{H}\log Z_{h}-\log Z=\Gamma.$ (68)
If heads produce non-proportional logit vectors, then $\Gamma>0$ by the strict convexity of log-sum-exp aggregation.
@@
which violates semantic information conservation principle. The total loss of the components strictly exceeds the loss of the aggregate system, indicating an ungrounded reduction in uncertainty (excess confidence). ∎
@@ 7.4 The Hallucination Mechanics
That brings us to the conclusion that the attention mechanism violates the semantic information conservation by generating over-confident prediction of the next token. That over-confidence in response is the artifact of the softmax nonlinearity and the Jensen-gap quantified in Theorem 9.
We should notice that there is a difference between the mathematical overconfidence and factual correctness. The correct prediction may well be a lucky guess created by the aggregation process itself rather than the evidence-based logical proof provided by the components. But that confidence coming out of nothing is one the essential mathematical components of creativity, imagination or fabrication.
@@ 8 Discussion and Speculations
8.1 Case Study: Escaping Impossibility in Hybrid Architectures
The constructive consequence of the impossibility theorems is the shift from attempting to eliminate hallucinations to designing systems that manage the trade-offs in a principled manner. In this section, we analyze a real-world scenario one can think of, demonstrating how it relabels and relocates the impossibility rather than resolving it.
8.1.1 A Proposed Solution: The Decoupled RAG Architecture
Consider a common scenario of introducing RAG systems to improve the quality of question answering, studied e.g. in Edge2024FromLT ; Sun2024ReDeEPDH ; Xu2024RetrievalAugmentedGW ; Asai2023SelfRAGLT . It has been recognized that an LLM designed to provide advice based on a trusted corpus of documents not only hallucinates its recommendations but also hallucinates the source quotes meant to ground its claims. To combat this, a hybrid, multi-stage architecture may be proposed.
Stage 1 (Retrieval): An information retrieval module fetches verbatim fragments from the trusted source documents based on the user’s query.
Stage 2 (Presentation): The system presents raw, unaltered text fragments directly to the human user, without any LLM-based processing.
Stage 3 (Constrained Synthesis): After the user has seen the source material, an LLM is to synthesize the pre-vetted fragments from Stage 2 into a summary.
That architecture is indeed a reasonable attempt to engineer around the hallucination problem by creating a verifiable information trail. However, as we will see, it does not escape the impossibility.
8.1.2 Analysis via the Impossibility Framework
The decoupled RAG architecture does not create a single response generation mechanism. Instead, it reallocates the responsibility for meeting each property of idealized response across its components (or module).
Truthfulness:
To impose truthful representation of available information (to avoid misrepresentation), the system delegates the assessment of retrieved content to the human user. By presenting the raw source text, the architecture bypasses the LLM for the primary act of verification. However, there is no guarantee that all relevant source text has been retrieved given the provided query.
Relevant Knowledge Revelation:
The quality of the final output is bounded by the retriever’s ability to find all relevant information (or whatever is labeled as ground truth in a database or a dataset, putting their completeness and representativeness aside). The LLM itself is absolved of this task. But that means, it is the responsibility of the user to define a query precise enough to get what is relevant. Designing such a prompt or (No)SQL database query is a challenge in a general case of non-trivial questions, so we should expect some violations of the knowledge revelation principle.
Semantic Information Conservation and Knowledge-Constrained Optimality:
By design, the prompt constrains the LLM to the provided knowledge retrieved from a database and asks for an optimal (accessible) synthesis. As we have already seen, database queries returning raw information stored in memory satisfy the property of Semantic Information Conservation at the cost of creativity. So, we may be optimistic in that regard. However, now we call LLM to synthesize the retrieved information and that is just another example of a non-trivial query. The LLM can still introduce subtle misinterpretations or ungrounded causal links to improve the summary’s narrative flow, which is a manifestation of the Jensen Gap.
The architecture is a great example of principled trade-off management. However, rather than being solved or hacked, the impossibility has been moved in this interesting and rather representative scenario. Let us also point out that putting ourselves in the loop introduces our own cognitive biases we should be able to address properly Kahneman:2011fj ; thaler2015misbehaving ; ariely2008predictably . Let us take a look at that challenge as well.
8.1.3 The Final Stage of Inference: The User’s Cognitive Mechanism
One can quite justifiably argue that the analysis is incomplete without considering the final node in the information chain: the human mind.
Presenting information to human does not guarantee its assimilation. Our ability to act as a perfect verifier is constrained by our own knowledge base and cognitive limits. If the source texts are technically dense or require specialized domain expertise, the knowledge they contain may not be truly accessible to us, even when presented. In such cases, we may be unable to spot a subtle hallucination in the LLM’s compelling summary. We have recognized that well enough in the preceding sections. Our well-documented cognitive biases, such as confirmation bias, priming or motivated reasoning, suggest that the cognitive trajectory of least resistance is to accept the compelling and accessible narrative maximizing our uncertainty reduction experience.
8.1.4 Managing the Inevitability
The decoupled RAG architecture represents one of approaches to managing the inevitability of hallucination control. It does not constitute an escape from the impossibility theorem, but it succeeds in providing the user with the opportunity for verification. The impossibility is not resolved, but relocated to our judgment.
8.2 Hallucination, Imagination, Intelligence
@@
Hallucination and creativity are fundamentally the same phenomenon viewed through different lenses. When beneficial, we call it imagination, creative insight, emergent understanding, conceptual synthesis, or eureka moment. When harmful, we label it hallucination, confabulation, fabrication or misinformation.
@@
We may very well argue the mathematical structure is identical in both cases. It is encoded generation of outputs whose confidence exceeds what can be justified by available evidence. This suggest (machine) intelligence may have much to do with a controlled violation of information conservation principle in productive ways. Indeed, the famous results in reinforcement learning and studies of dopamine neurons confirm that fundamental role of making and rewarding good predictions. Intelligence comes when we learn how to make bold, forward looking and correct guesses, when we have capacity for imagination to dream about what may be and what may that all mean, as wall as for imagining what happened and what would have happened in the past. Hallucination or imagination is an unconstrained exploration of ideas necessary to prepare for what may come next bennett2023brief ; christian2021alignment ; Dennett1991-DENCE . And without that over-confident guesses, creativity, exploration, imagination, new ideas cannot find their way to the outside world.
8.3 The Outer Limits of Reason
@@
Just as Gödel’s theorems godel1931formal ; godel1932vollstandigkeit showed that sufficiently powerful mathematical systems must be incomplete, the impossibility theorems in this paper demonstrate that sufficiently capable inference systems are bounded in their reasoning. Both results arise from the tension between expressiveness and consistency. But that is not all. The impossibility of simultaneously achieving truthfulness, conservation, revelation, and optimality relates to other fundamental limitations in mathematics and physics.
Heisenberg’s Uncertainty heisenberg1927anschaulichen tells we cannot simultaneously know the exact position and momentum of an elementary particle, which is a special case of Fourier uncertainty known in signal processing. Arrow’s Impossibility arrow1950difficulty shows we cannot design perfect voting systems aggregating universal orderings of alternatives, such that meets reasonable conditions of nondictatorship, Pareto efficiency, independence of irrelevant alternatives (or joint rationality of choice). Turing turing1936computable and Church church1936unsolvable show the halting problem is unsolvable, there is no general algorithm that can determine whether a program will eventually halt (stop running) or run forever.
These result jointly describe intelligence as a mechanism dealing with the impossibilities and violating them given the outer limits of reason yanofsky2016outer .
8.4 Dennett’s Multiple Drafts Model
The impossibility theorem presented in this paper provides (quite unexpectedly to the author) a mathematical foundations for many of brilliant insights of Daniel Dennett’s formulated in the Multiple Drafts model of Consciousness Dennett1991-DENCE . Where Dennett proposed a qualitative philosophical framework, we can provide the formal mathematical structure revealing why such a model is not merely plausible but implementable and refutable in LLM-based (philosophical) laboratory.
8.4.1 Multiple Drafts as Competing Agents
In Dennett’s model, consciousness emerges from multiple parallel processes of interpretation occurring simultaneously across distributed systems. The auction of ideas we have studied in this paper provides the precise mathematical foundations for that process. Each agent represents an interpretative process, knowledge encodes draft’s content, strategy represents a bid for consciousness.
In our transformer analysis in Section 7, each attention head generating logit vector $\mathbf{l}^{(h)}$ writes its own draft of the next token. When Head 1 votes for fox while Head 2 votes for dog, we observe as competing narratives bidding for selection.
8.4.2 Formation of AI Consciousness
Dennett argues that drafts become conscious (we will speculate what that might mean in a moment) only when probed. The LLM inference and sampling from fine-tuned inferred distributions formalizes this process. That, in turn, explains Dennett’s concept of confabulation, or the brain’s tendency to fill in gaps when creating coherent narratives from contradictory drafts.
By Theorem 8 and 9, the Jensen gap quantifies the excess confidence in reports that cannot be justified by constituent drafts. Without that overconfidence a victorious idea that reached the external world (as an LLM response) could loose the fierce competition with other ideas and remain inaccessible or hidden in the knowledge space. That suggests that the very excess confidence might be the mathematical signature of consciousness formation. In Dennett’s model and words, the subjective feeling of certainty emerges from this aggregation-induced distortion of information set.
The rigorous and careful neuro-scientific study of the evolution of brain and intelligence, beautifully summarized by Max Bennet in bennett2023brief , shows that at some point in history we developed the ability to model others’ minds. That ability provided great advantages in collecting information for future use. Understanding what someone else may think allows us to formulate a better context-related question that results in a better answer. When we anticipate understanding of our spoken message, we can make the knowledge hidden in someone else’s brain much more accessible. Therefore, as pointed out by Dennett, it does seem plausible that someone did accidentally hear his or hers own question fine-tuned to get into other mind sometime in the past, and that triggered emergence of an answer in the very same brain. That answer then could have made knowledge hidden in the author’s brain accessible to the author that was previously unaware, or unconscious if its existence or possession. The hunger for uncertainty reduction was satisfied as a result of an inner dialog.
We may therefore speculate that sustainable stimulation of that cognitive hunger, or the feeling of the cognitive hunger itself, is one of the components of consciousness formation. In our mechanistic or pragmatic framework, that could be described as a (nonlinear) dynamical system governed by the violations of semantic information conservation principle, defining attractors of ideas and self-referential concepts.
Let us acknowledge that is a bold hypothesis to be falsified and confronted with existing models of consciousness (see e.g., Chalmers chalmers1995facing or Levine levine2001purple ) and neuro-scientific discoveries (see e.g. cfang2025human or cogitate2025adversarial ). However, many of the building blocks of the consciousness formation process described above are provided in this paper.
8.4.3 A Critique
Dennett’s model has been criticized by many, and it is beyond the scope of this work to present that critique in its full capacity. Instead, let us only refer to the two basic points and speculate how the availability of LLMs could help philosophers refute available theories.
First, Searle john1997mystery and others argue that Dennett eliminates consciousness rather than explaining it, arguing that the spread of ideas is not driven by blind random forces but requires intentionality. The compelling and misleading nature of LLM responses that are generated by random sampling of trained distributions challenge that line of thinking. This paper suggests that the experiments focusing on LLM-hallucination detection may prove useful here to measure or identify intentionality (once we can have its programmable definition).
Second, there has been a charge that the multiple drafts model makes no concrete, falsifiable predictions, rendering it difficult to verify or integrate with neuroscience. This paper provides a mathematical toolkit for experimental philosophy to make that and other models a testable scientific hypothesis in the LLM-based laboratory.
@@ 9 Summary
We have established that no large language model can simultaneously achieve perfect hallucination (or imagination) control. We showed that four essential properties—truthful knowledge representation, semantic information conservation, complete revelation of relevant knowledge, and optimal response generation—are mutually incompatible. This impossibility emerges from the mathematical structure of information aggregation itself, not from limitations in data, compute, or architecture.
We introduced a rigorous mathematical framework and grounded in it the general analysis of inference processes and their LLM implementations. At the heart of our mathematical framework lies the semantic information measure $\mu_{C}$ and the emergence operator $\mathcal{E}_{C}$. The earlier one is a context-dependent metric capturing how knowledge reduces uncertainty within computational bounds. The later one formalizes how reasoning makes latent knowledge explicit rather than creating it anew. Finally, we model LLM inference as an auction of ideas—a idealized marketplace where neural components compete by bidding with their partial knowledge, each trying to influence the final response.
The impossibility manifests through three complementary proofs, each illuminating different aspects of the fundamental constraint. Through game theory, we apply the Green-Laffont theorem to show the impossibility when knowledge components are independent. Through probability theory, we leverage Savage’s proper scoring rules to extend the result to general trained or fine-tuned probability distributions. And through direct analysis of transformer architectures, we demonstrate how log-sum-exp convexity in attention mechanisms creates measurable violations of information conservation. The last proof exploits Jensen inequality for convex mappings to define the precise mathematical quantity that measures how aggregation of ideas (auction bids) produces overconfident responses that the constituent components cannot justify with revealed knowledge.
These mathematical insights we have made yield intriguing philosophical implications. Hallucination and creativity (two sides of the same coin we call imagination) emerge as mathematically identical phenomena, distinguished only by our normative judgments. The impossibility result explores the outer limits of reason, identified by Gödel’s incompleteness in mathematics, Heisenberg’s uncertainty in physics, and Arrow’s impossibility in social choice. Also, our framework seems to provide mathematical foundations and experimental setting for Dennett’s Multiple Drafts model of consciousness, suggesting that the measurable and observable excess confidence of LLM responses might be one of signatures of conscious processing of a primitive form.
Our goal was also to develop a theoretical framework that may open interesting avenues for future research. Some of ideas and bold speculations competing for attention can be mentioned here. In mathematical consciousness studies, we might formalize the notion of cognitive hunger as an attractor in knowledge space, sustained by continuous violations of information conservation. This could enable us to measure artificial consciousness signatures and test whether biological systems exhibit similar knowledge conservation violations. The framework suggests artificial consciousness might be studied as a nonlinear dynamical system whose attractors emerge from information creation events that our theorems describe. And so, the suggested implications may extend beyond language models to fundamental questions about intelligence itself.